OpenStudy (anonymous):

Using Gaussian elimination and back substitution solve the following system of equation

OpenStudy (anonymous):

x-4y-z=-13 -2x+3y+3z=2 3x-5z=17 9x+2y=-1

OpenStudy (anonymous):

I have a 4th row... does that mean that it's no solvable

OpenStudy (anonymous):



OpenStudy (anonymous):

Does that mean I ignore the last part and just find x y and z??

OpenStudy (amistre64):

in other words, you have 4 planes that might intersect at a common line or point, right?

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

sorry, i assumed you were familiar with the question you posted. im trying to get a sense of what this system of equations represents is all

OpenStudy (amistre64):

rref{{1,-4,-1,13},{-2,3,3,-2},{3,0,-5,-17},{9,2,0,1}} the wolf made nice work of that for me .... since the last row is all zeros, and 0 always equals zero, then the common intersection seems to be a plane ...

OpenStudy (anonymous):

so I am right in sayingx=-1 y=4 and z=-4

OpenStudy (amistre64):

my gut says yes .... but now youve got me wondering

OpenStudy (amistre64):

the row of zeros simply means that one of these equations can be written as a multiple of the others

OpenStudy (amistre64):

lets test it out :)

OpenStudy (amistre64):

x-4y-z=-13 -1 4 -4 ------- -1-16+4 = -13 -2x+3y+3z=2 -1 4 -4 ----------- 2 +12 -12 = 2 3x-5z=17 -1 -4 ------ -3+20 = 17 9x+2y=-1 -1 4 ------- -9+8 = -1 they fit

OpenStudy (anonymous):

Ahhhh good! I was just caught a little off guard by the fourth row haha thank you!!

OpenStudy (amistre64):

i agree that was a little off putting :) but as long as the last row didnt end in a nonzero value, were good